OpenStudy (jhannybean):

Surface Integrals for fun. \[\int\int_S xyz dS\] \(S\) is the cone with parametric equations \(x=u\cos(v)~,~ y=u\sin(v)~,~z=u~,~0\le u \le 1~,~ 0\le v \le \pi/2\)

OpenStudy (jhannybean):

since they've given us \(x~,y~,z\), we could write our parametric equation for our surface \(\vec r(u,v)\) I'm guessing.

OpenStudy (jhannybean):

No you were right, why delete it?

ganeshie8 (ganeshie8):

one sec deleted by mistake

OpenStudy (jhannybean):

\[\vec r(u,v) =\langle u\cos(v)~,u\sin(v)~,~u\rangle\]

ganeshie8 (ganeshie8):

\[\iint_S f(x,y,z)~ \mathrm{d}S = \iint_D f(\vec{r}(u,v))~\| \vec{r}_u \times \vec{r}_v\|~\mathrm{d}A\]

OpenStudy (anonymous):

To check your answer, I got the integral evaluated to \[ \frac{1}{5 \sqrt{2}} \] unless I did some small computation error.

OpenStudy (jhannybean):

Yeah, it was a small computational error.

ganeshie8 (ganeshie8):

im getting the same..

OpenStudy (jhannybean):

\[\frac{1}{10}\sqrt{2}\]

OpenStudy (jhannybean):

\[\left|\vec r_u \times \vec r_v\right| = \left[\begin{matrix}\hat i & \hat j & \hat k \\ \cos(v) & \sin(v) & 1\\ -\sin(v) & \cos(v) & 0 \end{matrix}\right] \\ =\langle 0-\cos(v)~,~ -(0+\sin(v))~,~ \cos^2(v) +\sin^2(v)\rangle = \langle -\cos(v)~,~ -\sin(v)~,~ 1\rangle \]

OpenStudy (jhannybean):

Oh, forgot to take the magnitude.

ganeshie8 (ganeshie8):

\[\left|\vec r_u \times \vec r_v\right| = \left|\begin{matrix}\hat i & \hat j & \hat k \\ \cos(v) & \sin(v) & 1\\ -\color{red}{u}\sin(v) & \color{red}{u}\cos(v) & 0 \end{matrix}\right| \\ \]

OpenStudy (jhannybean):

Oh, because w.r.t v.

ganeshie8 (ganeshie8):

yes i got the same ! just a notation thing : use vertical bars for determinant. square brackets refer to a matrix, not determinant

ganeshie8 (ganeshie8):

matrix A : \[A = \begin{bmatrix}a&b\\c&d \end{bmatrix}\] determinant of matrix A : \[|A| = \begin{vmatrix}a&b\\c&d \end{vmatrix}\]

OpenStudy (jhannybean):

\[\left|\vec r_u \times \vec r_v\right| = \left|\begin{matrix}\hat i & \hat j & \hat k \\ \cos(v) & \sin(v) & 1\\ -u\sin(v) & u\cos(v) & 0 \end{matrix}\right| \\=\langle 0-u\cos(v)~,~ -|(0+u\sin(v))~,~ u\cos^2(v) +u\sin^2(v)\rangle \\ = \langle -u\cos(v)~,~ -u\sin(v)~,~ u\rangle \\ =\left | \vec r_u \times \vec r_v \right| = \sqrt{u^2\cos^2(v) +u^2\sin^2(v)+u^2} =u\sqrt{2}\]

ganeshie8 (ganeshie8):

looks great!

OpenStudy (jhannybean):

\[\iint_S f(x,y,z)~ \mathrm{d}S = \iint_D f(\vec{r}(u,v))~\| \vec{r}_u \times \vec{r}_v\|~\mathrm{d}A\] \[x=u\cos(v)~,~y=u\sin(v)~,~z=u\]\[\iint_D (u\cos(v))(u\sin(v))(u)\cdot u\sqrt{2}dudv = \sqrt{2}\iint_D (u^3\cos(v)\sin(v))dudv \]

ganeshie8 (ganeshie8):

you should get u^4 in integrand..

ganeshie8 (ganeshie8):

\[\iint_D (\color{red}{u}\cos(v))(\color{red}{u}\sin(v))(\color{red}{u})\cdot \color{red}{u}\sqrt{2}dudv = \sqrt{2}\iint_D (u^{\color{red}{4}}\cos(v)\sin(v))dudv\]

OpenStudy (jhannybean):

Oh I missed a u :(

OpenStudy (jhannybean):

miscounted -.-

ganeshie8 (ganeshie8):

bounds and the differential are not agreeing.. check once

OpenStudy (jhannybean):

\[\sqrt{2} \int_0^1\int_0^{\pi/2} \left(u^4 \cos(v)\sin(v)\right)dvdu\]

ganeshie8 (ganeshie8):

wolfram it :)

OpenStudy (jhannybean):

:P

OpenStudy (jhannybean):

I can split the integrals, then on the v integral just use a u-sub, right?

OpenStudy (jhannybean):

\[\sqrt{2}\left[\int_0^1 u^4du \int_0^{\pi/2} \cos(v)\sin(v)dv\right]\]

OpenStudy (jhannybean):

let \(u=\sin(v)~,~ du = \cos(v)dv\)\[\sqrt{2} \left[\left[\frac{1}{5}u^5\right]_0^1 \cdot \left[\int_0^1 udu\right]\right] = \sqrt{2}\left[\frac{1}{5}\cdot \left. \left(\frac{1}{2}u^2\right|_0^1 \right) \right] =\sqrt{2} \cdot\frac{1}{5} \cdot \frac{1}{2} =\frac{1}{10}\sqrt{2} \]

OpenStudy (jhannybean):

There we go ) Just a fun, refreshing surface integral before I headed to sleep :P

ganeshie8 (ganeshie8):

Wow! you have evaluated manually and managed to arrive at the correct answer! good job :)

OpenStudy (jhannybean):

Haha, thanks!! It's good practice.